OpenStudy (anonymous):

Which of the below systems of equations matches the following graph? x – y = 1 3x – 2y = 8 x – y = –1 3x + 2y = –8 x + y = 1 3x – 2y = 8 x + y = –1 3x + 2y = –8

OpenStudy (anonymous):

The graph is attached below

OpenStudy (anonymous):

ur graph is too small to see and u will just have to solve the equations first.

OpenStudy (anonymous):

JayDS if u click the link it becomes bigger :)

OpenStudy (anonymous):

for the first one of x-y=1 3x-2y=8 we can firstly rearrange x-y=1 to be x=1 + y so y=x-1 then we sub y=x-1 into the second equation of 3x-2y=8 3x-2(x-1)=8 3x-2x+2=8 x=8-2 x=6 now sub x=6 back into either equations, lets choose y=x-1 so y=6-1 y=5 so the point of intersection here is (6,5)

OpenStudy (anonymous):

and sorry, what you said for your graph does not work.

OpenStudy (anonymous):

oh okay im sorry

OpenStudy (anonymous):

Do you understand what I did for the first example?

OpenStudy (anonymous):

yea :)

OpenStudy (anonymous):

ok now, do it for the rest.

OpenStudy (anonymous):

umm...i think the answer is a..........??

OpenStudy (anonymous):

is it? :D

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

did u actually solve it?

OpenStudy (anonymous):

you can never just guess.

OpenStudy (anonymous):

i did solve it....didnt just guess

OpenStudy (anonymous):

the point of intersection is (-2,-1) on the graph so you have to solve all the equations to find their point of intersection that matches that.

OpenStudy (anonymous):

It is clearly not the first one because as I solved above the point of intersection is (6,5)

OpenStudy (anonymous):

so go through your working out again and re-do it if you have to.

OpenStudy (anonymous):

I have to go but just use your calculator to check afterwards if you are correct or not by plotting both equations each time and looking at the graph.